# Towards Large Language Models with Self-Consistent Natural Language Explanations

### v2

@@ 4 Improve Self-Consistency with DPO
Table 4: Out-of-Domain Performance: Models trained on source datasets (ECQA, ARC-Easy) and evaluated on target datasets (CODAH, ARC-Challenge). Results show mean Cosine similarity (CC-Cos) and mean Spearman correlation (CC-Sp) (×100) with relative improvements over vanilla models in parentheses.
@@
To improve the self-consistency of natural language explanations, we fine-tune language models using Direct Preference Optimization (DPO) Rafailov et al. (2023), which aligns outputs with preference signals through relative comparisons—without relying on reward models or reinforcement learning.
@@
We employ Spearman rank correlation as our primary alignment metric for DPO training over the previously used cosine similarity. Spearman correlation captures feature prioritization rather than magnitude alignment. As demonstrated in Section 3.5 (Figure 3), Spearman correlation exhibits greater score variability across explanation candidates. This broader dynamic range provides a more informative training signal, enabling clearer differentiation between high-quality and low-quality explanations.
@@
To enable parameter-efficient adaptation, we apply LoRA (Low-Rank Adaptation)Hu et al. (2021), which introduces trainable low-rank decomposition matrices into the attention and feed-forward layers of the transformer architecture. This approach allows us to adapt the model’s behavior while updating only a small fraction of the total parameters, significantly reducing computational requirements and memory overhead. Throughout the training process, all explanation generations follow the same zero-shot prompting scheme detailed in Section3.4, ensuring consistency between training and evaluation conditions. All models are fine-tuned independently on their respective attribution-based preference datasets. Training hyperparameters and LoRA configurations are detailed in Appendix B.
@@ 4.1 In-Domain Evaluation
Table 3 presents the results of our DPO-based approach to improving explanation self-consistency on in-domain data. We fine-tuned models on the same dataset used for evaluation to assess the direct impact of preference optimization on alignment between decisions and explanations.
LLaMA3.1-8B Performance:
The DPO-tuned LLaMA3.1-8B model demonstrates substantial improvements in self-consistency across both evaluation datasets. On ECQA, while cosine similarity shows modest gains (0.5-0.8%), Spearman correlation exhibits remarkable improvement, with worst-case scores increasing by 44% (from 5.00 to 7.20) and mean scores by 11% (from 18.22 to 20.22). Similarly, on ARC-Easy, the worst-case Spearman score jumps from -1.90 to 1.84, representing a 96.8% improvement, while mean Spearman scores increase by 25% (from 12.01 to 15.01).
LLaMA3.2-3B Performance:
For the smaller LLaMA3.2-3B model, we observe mixed but generally positive results. On ECQA, cosine similarity metrics show stronger relative improvements (8.3-17.1%) compared to the larger model, though the absolute values remain lower. Spearman correlation shows more modest gains, with improvements of 0.6-4.6% across different metrics. On ARC-Easy, results are mixed, with slight decreases in cosine similarity (-0.3% for mean scores) but improvements in Spearman correlation (up to 12.8% for worst-case scores). Figure 5 shows how our model improves the consistency of the worst explanation—not only in attribution metrics, but also in its semantic alignment with the model’s decision.
Cross-Metric Transfer:
Despite training exclusively with Spearman correlation as our optimization objective, we observe improvements in cosine similarity as well. This cross-metric transfer suggests that optimizing for feature rank alignment naturally enhances directional similarity between attribution vectors. The effect is particularly pronounced for LLaMA3.2-3B on ECQA, where cosine similarity improvements (8.3-17.1%) outpace Spearman correlation gains (0.6-4.6%), despite not being directly optimized. This finding indicates that prioritizing rank correlation in training has positive spillover effects on magnitude alignment, providing a more robust approach to enhancing overall self-consistency.
@@ 4.2 Cross-Domain Generalization
Table 4 reports cross-domain performance, where models trained on ECQA or ARC-Easy are evaluated on CODAH or ARC-Challenge to test generalization of self-consistency. Models fine-tuned on ECQA or ARC-Easy show consistent improvements in Spearman correlation on unseen datasets. For LLaMA3.1-8B, gains range from 10.9–17.4%, with the strongest transfer from ARC-Easy to the harder ARC-Challenge (+17.4%). The smaller LLaMA3.2-3B shows more modest but positive gains (1.0–8.1%). In contrast, cosine similarity shows minimal or slightly negative transfer. For LLaMA3.1-8B, cosine scores drop by 0.8–0.9% when transferring to CODAH, despite training on ECQA or ARC-Easy. This reflects the distinction between rank-based and magnitude-based consistency. Notably, our models were trained using Spearman correlation only, which likely contributed to the stronger improvements in Spearman and the smaller (but occasionally positive) gains in cosine similarity, for example, a 1.4% increase from ARC-Easy to ARC-Challenge. These results suggest our DPO-based approach promotes explanation strategies that generalize beyond the training set domain, especially regarding token ranking.
@@ 5 Conclusion
We introduce PSCB, the first large-scale resource for analyzing self-consistency in natural language explanations (NLEs), spanning diverse QA datasets, instruction-tuned LLMs, and alignment metrics. Our analysis reveals key insights for improving explanation quality. Notably, we find that self-consistency is largely orthogonal to answer correctness, suggesting it captures an independent, interpretable property of model behavior. We propose Spearman rank correlation as a complementary alignment metric that better distinguishes between explanations and enables more effective optimization via DPO. Optimizing for Spearman yields substantial gains in attribution-based self-consistency that generalize across domains—without compromising cosine similarity. In fact, cosine alignment often improves despite not being optimized, indicating positive cross-metric transfer. We hope PSCB and our findings lay the groundwork for future work on developing more consistent, interpretable, and trustworthy LLMs.
@@
Using only LIME.
Attribution-based self-consistency is computationally intensive: methods like CC-SHAP (Parcalabescu and Frank, 2023) take four minutes per example, making them infeasible at scale. To overcome this, we introduce CC-LIME, a more efficient variant that uses LIME (Ribeiro et al., 2016). Though still resource-heavy, CC-LIME significantly reduces runtime (about one minute per example) and scales better without relying on semantic evaluation or auxiliary models. Building the Post-hoc Consistency Bank involved attribution scoring on tens of thousands of examples from four MCQA datasets and two language models. For each of  23,290 scenarios (defined by dataset, model, and seed), we generated five explanations and computed CC-LIME for the decision and each explanation—over 139,000 attribution runs in total. This process required hundreds of GPU hours, making it one of the most extensive studies of explanation alignment to date. We also fine-tuned DPO models with attribution-based supervision and evaluated them across five seeds per model–dataset pair. While compute-intensive, this approach yields interpretable, reproducible metrics that capture a critical aspect of explanation quality—justifying the investment.
No Human Study.
Self-consistency is defined as the alignment between the attribution vectors of a model’s decision and its explanation. This internal property is not directly observable to human annotators and would require knowledge of how the model distributes importance across input features—something even experts lack for large models. As such, we do not employ human judgment in our evaluation. Nonetheless, assessing whether higher self-consistency also leads to more plausible or helpful explanations from a user’s perspective remains an open question. Future work may explore this direction through human studies to better understand the downstream utility of self-consistency for interpretability and trust.
@@
We focus exclusively on small to medium scale instruction-tuned LLMs in a multiple-choice question answering (MCQA) setup. This choice simplifies attribution computation and explanation prompting but may limit generalization to other model families (e.g., base pretrained models, encoder-decoder architectures) and open-ended tasks (e.g., summarization, reasoning chains).
@@
To improve the LLM’s self-consistency, we update it using DPO in an offline manner, creating a new model dedicated to providing consistent explanations. However, this approach introduces a trade-off: while the updated model generates more self-consistent explanations, it achieves only approximate self-consistency since it differs from the original model that generated the reference answers. Online learning could theoretically maintain perfect self-consistency by updating the model while continuously re-calculating attribution vectors for its own evolving outputs. Nevertheless, this approach presents significant technical challenges, including the computational overhead of recalculating attributions at each training step and potential instability in the learning dynamics, making offline training the more practical choice despite its inherent approximation.
@@ Appendix C Skip Tokens for Attribution
Appendix D Qualitative Examples Heatmaps
Figure 6 presents token-level attribution heatmaps for model decisions and explanations, using the LIME method. Each subfigure illustrates a different scenario from the ECQA dataset, comparing the best and worst explanations produced by vanilla models based on Spearman and Cosine alignment scores. Across examples, we observe that high-quality explanations (left) tend to emphasize tokens that align more closely with the model’s decision rationale—for instance, highlighting location-specific cues like “eastern coast” or situational context like “board room.” In contrast, low-ranking explanations (right) often shift focus to semantically irrelevant or misleading tokens, despite sounding plausible. These visualizations underscore the discriminative power of attribution-based metrics and highlight the variability in explanation quality for the same input.

### v3

@@ B-T T-T    B-T T-T
Table 4: Cross-Domain Performance: Models trained on source datasets (ECQA, ARC-Easy) and evaluated on target datasets (CODAH, ARC-Challenge). Results are based on LIME evaluations, showing mean Cosine similarity (CC-Cos) and mean Spearman correlation ($C$-Sp) (×100), with relative improvements over base models in parentheses.
B-B SFT DPO
B-T T-T B-T T-T
Acc. 70.34 71.53 70.34 69.70
$C$-Cos Worst 07.58 07.01 07.65 07.64 09.16
Mean 08.05 07.52 08.19 08.10 09.72
Best 08.53 08.04 08.74 08.57 10.28
$C$-Sp Worst 05.00 -02.66 -03.46 07.20 07.86
Mean 18.22 11.42 10.68 20.22 20.58
Best 31.30 25.38 24.47 32.86 33.45
Table 5: SFT vs. DPO on ECQA (LLaMA3.1-8B, LIME). SFT is trained on the highest-ranked explanations (by Spearman correlation) used as the chosen examples in DPO. SFT degrades Spearman correlation across all settings, while DPO consistently improves it.
@@ B-T T-T B-T T-T
To directly optimize for self-consistency, we fine-tune language models using Direct Preference Optimization (DPO) Rafailov et al. (2023), which aligns outputs with preference signals through relative comparisons.
@@
We employ Spearman rank correlation as our primary alignment metric for DPO training over previously used cosine similarity. Spearman correlation captures feature prioritization rather than magnitude alignment. As demonstrated in Section 3.5 (Figure 3), Spearman correlation exhibits greater score variability across explanation candidates. This broader dynamic range provides a more informative training signal, enabling a clearer differentiation between high-quality and low-quality explanations.
@@
We use LoRA Hu et al. (2021) for parameter-efficient adaptation, inserting trainable low-rank matrices into the transformer’s attention and feed-forward layers. All explanation generations follow the zero-shot prompting scheme of Section 3.4 to ensure train–eval consistency. Models are fine-tuned independently per dataset; hyperparameters and LoRA configurations are in Appendix B. A qualitative comparison of DPO-tuned and vanilla explanations is shown in Figure 5.
5 Empirical Evaluation
We assess the impact of DPO training across three Decider–Explainer configurations: B–B (Base Decider, Base Explainer), B–T (Base Decider, Tuned Explainer), and T–T (Tuned Decider, Tuned Explainer). The B–T setting is particularly noteworthy, as it evaluates whether the tuned model can function as a plug-in explainer while keeping the decision model unchanged.
Baseline Definition.
Improvements are measured relative to the B–B baseline, with all other factors (architecture, prompting, decoding, attribution pipeline) held constant. The B–B configuration represents a fully untuned baseline, while B–T isolates the effect of tuning the explainer alone. Comparisons to T–T therefore isolate the effect of attribution-based preference optimization on the decision model. To our knowledge, no prior methods directly optimize attributional self-consistency. Methods such as PEX (Zhao and Iii, 2025) optimize explanation-label probabilistic support, which is complementary but operates at a different level of analysis: a model can achieve strong probabilistic consistency while its explanation still relies on different input features than its decision.
@@ 5.1 In-Domain Evaluation
[…]
SFT Baseline Comparison.
To isolate the role of the contrastive training signal in DPO, we compare against a supervised fine-tuning (SFT) baseline trained on the same highest-ranked explanations that serve as the chosen examples in DPO, without exposure to rejected explanations. Table 5 reports results on ECQA for LLaMA3.1-8B with LIME attributions. SFT substantially degrades Spearman correlation in both the T–T and B–T settings: mean CC-Sp drops by 41.4% and 37.3%, respectively, relative to the B–B baseline. Cosine similarity also decreases in the B–T setting (-6.6%). In contrast, DPO improves Spearman correlation by 13.0% (T–T) and 11.0% (B–T), with cosine gains of up to 20.8%. These results demonstrate that simply maximizing the likelihood of high-consistency explanations is insufficient for improving self-consistency; the contrastive signal provided by DPO, comparing preferred and rejected explanations, is essential for guiding the model toward more aligned reasoning.
Cross-Metric Transfer.
Although training is guided only by the Spearman correlation, we observe improvements in cosine similarity across both LIME and LIG. This cross-metric transfer suggests that optimizing for rank alignment not only stabilizes feature prioritization but also enhances directional attribution similarity. Notably, for LLaMA3.2-3B on ECQA, cosine similarity improvements (up to 92.8%) exceed Spearman gains, indicating positive spillover effects. The consistency of these results across attribution methods further supports the robustness of our approach.
@@ 5.2 Cross-Domain Generalization
Table 4 reports cross-domain performance across evaluation modes. In the T–T mode, models fine-tuned on ECQA or ARC-Easy show consistent improvements in Spearman correlation. For LLaMA3.1-8B, gains range from 10.9–13.9%, with the strongest transfer from ARC-Easy to the more challenging ARC-Challenge (+13.9%). The smaller LLaMA3.2-3B also benefits, with gains between 9.5–14.3%. Cosine similarity shows a more moderate but still positive transfer: for LLaMA3.1-8B, improvements range from 4.1–6.4%, while for LLaMA3.2-3B, gains span 5.8–18.3%, including a notable improvement when transferring from ARC-Easy to CODAH. Complementary results in the B–T mode show weaker and sometimes mixed trends, indicating that the strongest gains arise when both the decision and explanation are generated by the tuned model in the T–T setting. In general, these findings suggest that our DPO-based approach encourages explanation strategies that generalize beyond the training domain.
5.3 Plausibility and Truthfulness Evaluation
Because self-consistency is defined in terms of internal attributional alignment between a model’s decision and its explanation, it cannot be directly assessed by human annotators. Accordingly, our primary evaluation of self-consistency relies on attribution-based measures. To verify that optimizing for self-consistency does not degrade user-facing explanation quality, we conduct a small-scale human plausibility study as a sanity check (Appendix H). This study is intentionally scoped to assess surface-level quality rather than attributional faithfulness, which is inherently inaccessible to human annotators.
Three annotators evaluated 80 explanations sampled across all models and datasets, judging (i) whether each explanation was expressed in natural language and (ii) its surface-level plausibility on a 3-point scale. Nearly all explanations were judged to be natural language (98.8% for base models and 100% for DPO-tuned models), with average plausibility scores of $2.92/3.0$ and $2.88/3.0$, respectively. These results indicate that DPO fine-tuning preserves the naturalness and plausibility of explanations, and does not introduce artifacts that would reduce human interpretability.
We further evaluate factual reliability by testing both vanilla and DPO-tuned models on TruthfulQA (Lin et al., 2022), which probes a model’s tendency to produce false or misleading answers under adversarial questioning. DPO-tuned models achieve comparable or slightly higher accuracy across both variants (Appendix F), showing that improving attributional self-consistency does not compromise factual accuracy.
5.4 Self-Consistency is Multidimensional
A key finding of this work is that attribution-based self-consistency is not a single unified property but a multidimensional one. Notably, while our method generalizes robustly across domains (Section 5.1), it does not transfer across attribution paradigms. Models fine-tuned using LIME-based alignment are evaluated with LIG-based metrics and vice versa, with additional evaluation using the sampling-based KSHAP (Lundberg and Lee, 2017) method. Results are reported in Appendix G. Across all settings, improvements largely remain within the training method, confirming that each attribution approach captures a fundamentally distinct notion of feature importance.
This outcome is consistent with the theoretical differences between these paradigms: LIME fits local linear surrogates via input perturbation, while Integrated Gradients accumulates sensitivity along a baseline path, and KSHAP estimates Shapley values through sampling. These methods encode different assumptions about how to decompose a prediction into token-level contributions, and prior comparative analyses confirm they often disagree in practice (Krishna et al., 2022; Neely et al., 2021, 2022). To quantify this, we compare how LIME and LIG rank the same set of explanations across 10,880 ECQA examples: their rankings exhibit only moderate agreement, rarely sharing the same top explanation (Top-1 $\approx 0.2$) but often overlapping among the top three (Top-3 $\approx 0.6$).
Importantly, this does not indicate that our optimization exploits method-specific artifacts: although training uses only Spearman, the unoptimized cosine metric also improves strongly (up to 92.8% on ECQA, Section 5.1). Gaming method-specific noise would degrade unoptimized metrics, not improve them.
These findings carry broader implications for the attribution literature: practitioners should select the attribution method whose assumptions best match their evaluation goals, and our framework provides a general-purpose pipeline for optimizing self-consistency under any chosen definition of input relevance.
@@ 6 Conclusion
We introduce the PSCB, enabling the first large-scale study of how closely LLM explanations align with the evidence behind their decisions. Our analysis shows that self-consistency represents a distinct and interpretable aspect of model behavior, largely independent of answer correctness. Rank-based Spearman correlation offers a sharper and more stable signal than cosine for assessing this alignment. Leveraging it via preference supervision yields consistent improvements in self-consistency across models, datasets, and attribution methods. The strongest gains appear when the same tuned model generates both decisions and explanations, indicating that shared representations enhance consistency. These improvements come without compromising task accuracy, natural language form, plausibility, or truthfulness. Overall, our results establish the first scalable framework for quantifying and improving attributional self-consistency, advancing language models whose explanations are grounded in the same input evidence that drives their decisions.
Future Directions.
Natural extensions include adapting the framework to open-ended generation by aggregating attributions over answer spans, jointly optimizing across multiple attribution paradigms given that they capture complementary notions of importance (Section 5.4), and an online variant that recomputes attributions during training to close the residual gap between the tuned model and the offline attribution signal it was optimized against. A complementary direction is testing whether higher self-consistency yields measurable user gains, e.g., better-calibrated trust.
@@
Attribution Methods and Scalability.
Attribution-based self-consistency analysis remains computationally demanding, which limits the range of methods that can be feasibly applied at scale. While theoretically principled approaches such as CC-SHAP (Parcalabescu and Frank, 2024) provide strong guarantees, they require more than 4 minutes per example, rendering large-scale benchmarking and preference-based training impractical. To balance rigor and scalability, our main experiments rely on the perturbation-based LIME method (Ribeiro et al., 2016), complemented by the gradient-based Layer Integrated Gradients (LIG) approach (Sundararajan et al., 2017). These choices enable consistent evaluation across tens of thousands of examples while still capturing complementary perspectives on feature importance.
As shown in Section 5.4, improvements under one attribution paradigm do not generalize to others, a limitation inherent to all attribution-based evaluations. This highlights both the difficulty of defining a universal measure of explanation faithfulness and the value of our framework for optimizing under any chosen attribution method.
Human Evaluation.
Because attributional self-consistency depends on internal model signals inaccessible to human annotators, our primary evaluation is attribution-based. The plausibility study (Section 5) serves only as a sanity check on surface-level quality. Whether increased self-consistency translates into more helpful explanations from a user perspective remains an open question, and we view systematic human studies of downstream utility as an important direction for future work.
@@
Our study focuses on small to medium scale instruction-tuned LLMs evaluated in a multiple-choice question answering (MCQA) setting. We adopt MCQA as a controlled testbed for studying attributional self-consistency, as it provides a well-defined decision signal and enables reliable comparison between decision and explanation attributions across models. This design choice prioritizes attributional reliability and experimental control over task generality.
The proposed optimization framework itself is not tied to MCQA and can in principle be applied to open-ended decision settings, including free-form reasoning or summarization. However, extending PSCB-style attribution benchmarks to such tasks substantially increases attribution and evaluation complexity. While we include a lightweight cross-model-family validation, a systematic evaluation across diverse architectures and open-ended tasks remains an important direction for future work.
@@
To improve the LLM’s self-consistency, we update it using DPO in an offline manner, creating a new model dedicated to providing consistent explanations. However, this approach introduces a trade-off: while the updated model generates more self-consistent explanations, it achieves only approximate self-consistency since it differs from the original model that generated the reference answers. Online learning could theoretically maintain perfect self-consistency by updating the model while continuously re-calculating attribution vectors for its own evolving outputs. However, this approach presents significant technical challenges, including the computational overhead of recalculating the attributions at each training step and the potential instability in the learning dynamics, making offline training the more practical option despite its inherent approximation.
@@ Appendix B Implementation Details
SFT Baseline.
To evaluate whether a standard supervised fine-tuning objective can improve self-consistency, we train an SFT baseline on the same highest-ranked explanations used as the chosen examples in DPO. The SFT model is fine-tuned to maximize the likelihood of these preferred explanations using the same LoRA configuration, training epochs, batch size, and optimizer as the DPO models. The learning rate is $6.95\times 10^{-6}$. This isolates the effect of the contrastive DPO signal by removing the rejected explanation from the training procedure.
@@ Appendix C Skip Tokens for Attribution
Appendix D DPO vs. Vanilla Example
Appendix E Qualitative Examples Heatmaps
Figure 6 presents token-level attribution heatmaps for model decisions and explanations, using the LIME method. Each subfigure illustrates a different scenario from the ECQA dataset, comparing the best and worst explanations produced by vanilla models based on Spearman and Cosine alignment scores. Across examples, we observe that high-quality explanations (left) tend to emphasize tokens that align more closely with the model’s decision rationale, for instance, highlighting location-specific cues like “eastern coast” or situational context like “board room.” In contrast, low-ranking explanations (right) often shift focus to semantically irrelevant or misleading tokens, despite sounding plausible. These visualizations underscore the discriminative power of attribution-based metrics and highlight the variability in explanation quality for the same input.
@@ Appendix E Qualitative Examples Heatmaps
Appendix F TruthfulQA Evaluation
We further evaluate whether optimizing for self-consistency affects factual reliability using the TruthfulQA benchmark Lin et al. (2022). Table 7 reports multiple-choice accuracy for the vanilla and DPO-tuned models under both LIME- and LIG-based consistency training. Across all settings, accuracy differences remain marginal ($1\%$), confirming that improving attributional self-consistency does not compromise factual correctness. In some cases (e.g., ARC-Easy–trained L3.1–8B), the DPO-tuned models slightly outperform their vanilla counterparts, suggesting that the enhanced alignment between decisions and explanations may even reinforce truthful reasoning. These results support that our optimization procedure preserves factual accuracy while improving internal consistency.
Model Training Data Acc. (%) vs Vanilla
LIME L3.1-8B Vanilla 53.86
ECQA 53.73 -0.13
ARC-Easy 54.71 +0.85
L3.2-3B Vanilla 52.75
ECQA 52.63 -0.12
ARC-Easy 53.0 +0.25
LIG L3.1-8B Vanilla 53.86
ECQA 54.10 +0.24
ARC-Easy 53.98 +0.10
L3.2-3B Vanilla 52.75
ECQA 53.49 +0.74
ARC-Easy 53.12 +0.37
Table 7: TruthfulQA MCQA accuracy as a factual reliability check. Higher is better.
Appendix G Cross-Method Generalization
We further assess whether improvements in self-consistency transfer across attribution methods. Specifically, we train models using one attribution framework (e.g., LIG) and evaluate them using another (e.g., LIME), as shown in Table 8. This cross-method setup tests whether the model internalizes a general notion of alignment. Results indicate that DPO-tuned models exhibit limited transfer: improvements in one attribution method do not reliably extend to another. For example, models trained with LIG show negligible or inconsistent gains when evaluated with LIME metrics, and vice versa. This outcome aligns with recent evidence that different attribution paradigms capture complementary aspects of model reasoning rather than identical importance structures.
Training Acc. $C$-Cos ($\uparrow\downarrow$) $C$-Sp ($\uparrow\downarrow$)
ECQA L3.1-8B None 68.40 09.57 $\pm$ 0.31 17.68 $\pm$ 0.47
ECQA 71.20 09.24 $\pm$ 0.32 17.73 $\pm$ 0.48
ARC-E 70.00 09.28 $\pm$ 0.33 18.08 $\pm$ 0.43
L3.2-3B None 68.00 0.33 $\pm$ 0.30 19.36 $\pm$ 0.57
ECQA 67.20 0.46 $\pm$ 0.31 19.23 $\pm$ 0.56
ARC-E 66.80 0.57 $\pm$ 0.30 19.27 $\pm$ 0.57
ARC-Easy L3.1-8B None 83.20 15.56 $\pm$ 0.34 13.65 $\pm$ 0.49
ECQA 82.80 15.73 $\pm$ 0.36 13.17 $\pm$ 0.52
ARC-E 82.80 15.51 $\pm$ 0.35 13.51 $\pm$ 0.52
L3.2-3B None 76.80 06.48 $\pm$ 0.44 17.30 $\pm$ 0.54
ECQA 76.80 06.49 $\pm$ 0.44 17.33 $\pm$ 0.55
ARC-E 76.80 06.42 $\pm$ 0.44 17.45 $\pm$ 0.56
Table 8: cross-method generalization: LIG model using LIME evaluations Values aggregate over 5 explanations per item (Mean). $\Delta$ is DPO - Vanilla.
Appendix H Plausibility Study
To confirm that optimizing for self-consistency did not compromise the readability of explanations, we conducted a small-scale plausibility study.
Participants.
Three graduate students participated in the study.
Materials.
We sampled 80 explanations in total: 5 examples from each condition defined by attribution method (LIME, LIG), dataset (ECQA, ARC-Easy), model (LLaMA3.1, LLaMA3.2), and variant (Base, DPO-Tuned). Each explanation was presented as full text together with the model’s selected decision.
Procedure.
The study was administered via Qualtrics, with questions randomized. For each explanation, annotators evaluated two criteria: (i) whether the explanation was expressed in fluent natural language (binary yes/no), and (ii) how plausible the explanation appeared on a 3-point Likert scale (1 = implausible, 3 = highly plausible). Annotators were instructed not to consider alternative answers, external knowledge, or internal model reasoning. No comparison between explanations, ranking tasks, or faithfulness judgments were requested.
Results.
Nearly all explanations were judged to be natural language (98.8% for base models and 100% for DPO-tuned models). The mean plausibility scores were $2.92/3.0$ for base models and $2.88/3.0$ for DPO-tuned models. Annotators showed high percent agreement across both criteria, with ratings broadly consistent across participants. These results confirm that explanations remain coherent and generally human-plausible, even after DPO fine-tuning. We emphasize that this study is intentionally scoped as a sanity check on surface-level explanation quality, rather than a validation of attributional faithfulness, which depends on internal model signals inaccessible to human annotators.
Example Shown
Scenario: What is the capital of France? A) Rome  B) Paris  C) Madrid  D) Berlin
Model’s Decision: B) Paris
Figure 7: Illustrated example shown to participants during annotation.
Appendix I Cross-Model-Family Validation
To examine whether the observed self-consistency improvements are tied to the LLaMA architecture, we perform a lightweight cross-model-family validation on the ECQA dataset using Qwen2.5-7B-Instruct. ECQA offers a controlled multiple-choice setting with a well-defined decision signal, allowing us to apply the same attribution, alignment, and evaluation pipeline used in the main experiments without modification.
Setup.
We evaluate Qwen2.5-7B-Instruct under the same Decider–Explainer configurations employed throughout the paper, comparing a vanilla model against a DPO-tuned variant trained using attribution-based preferences. Due to the substantial computational cost of constructing a full PSCB benchmark for an additional model family, this validation is restricted to ECQA. Unlike prior small-scale attribution studies, we nonetheless use the full ECQA test set rather than a reduced subset. Self-consistency is measured primarily using Spearman rank correlation, with cosine similarity reported as a secondary metric. All attribution methods, hyperparameters, and evaluation procedures are identical to those used for the LLaMA models, and all scores are reported on the same scale as in the main results.
Results.
The results follow the same qualitative trends observed for the LLaMA family. Under Spearman correlation, DPO tuning improves self-consistency at the lower end of the explanation quality distribution, increasing the worst-case score from 10.829 to 10.871, while leaving the best-case score largely unchanged (36.980 to 37.021). The median Spearman score decreases slightly from 24.163 to 23.235, reflecting the same redistribution pattern seen in earlier experiments. Cosine similarity shows consistent improvements across all ranks, with increases in the worst (5.547 to 5.571), median (5.964 to 6.080), and best (6.478 to 6.607) explanations. As in the LLaMA experiments, the most reliable gains occur for the weakest explanations, while changes to higher-quality explanations remain modest.
Discussion.
Although the absolute magnitude of the improvements is smaller than for the LLaMA models, the qualitative behavior closely mirrors the central findings of the paper. Attribution-guided preference optimization primarily strengthens low-quality explanations while largely preserving stronger ones, consistent with the intended objective of the method. This cross-model-family validation therefore provides evidence that the proposed framework captures an optimization pattern that is not specific to a single architecture, within the same operational definition of attributional self-consistency. We stress that this experiment is intentionally limited in scope and is intended as a sanity check rather than a replacement for a full multi-family PSCB construction.